OpenStudy (anonymous):

Determine whether the series is convergent or divergent: summation n=1 to infinity 1/(n^(3)-n) I walked through using a similar example as a guide and found the series to be convergent by use of the integral test. Just wanted to make sure this is correct because I only have one submission on this problem

OpenStudy (rational):

are you sure the index starts at 1 and not 2 ?

OpenStudy (rational):

because plugging n=1 gives 1/(1^3-1) = 1/0 as first term

OpenStudy (anonymous):

Yes, that's why I'm getting hung up. I've done the one with n=2 before, but not n=1

OpenStudy (rational):

wolfram says sum doesn't exist http://www.wolframalpha.com/input/?i=summation+n%3D1+to+infinity+1%2F%28n%5E%283%29-n%29

OpenStudy (rational):

it would be a different story if the index starts from 2

OpenStudy (jhannybean):

\[\sum_{i=1}^{\infty}\frac{1}{n^3-n}\implies \sum_{i=1}^\infty\left(\frac{1}{n^3}-\frac{1}{n}\right) \]

OpenStudy (anonymous):

if n=2, it would be conv - I just didn't know how to handle the n=1

OpenStudy (rational):

i would listen to wolfram

OpenStudy (jhannybean):

You went through these? http://www.math.utah.edu/~tskorc/1220/convergence_tests.pdf

OpenStudy (jhannybean):

Looks kind of like p-series

OpenStudy (anonymous):

We've gotten to limit comparison in class, but this problem in in the area where we're doing integral tests

OpenStudy (jhannybean):

And you haven't learned p-series? That's probably the easiest convergent series tests out of all of them

OpenStudy (jhannybean):

You're just comparing the powers.

OpenStudy (anonymous):

We got to p_series, because limit comparison is when p-series fails

OpenStudy (jhannybean):

Try p-series.

OpenStudy (anonymous):

Using p-series, it would Conv. When I used partial fractions and the integral test, I got 1/2 ln(1) =0. What I don't get is the n=1 because it starts the series with an undefined value.

OpenStudy (rational):

technically speaking the sum doesn't exists because that undefined term ruins everything

OpenStudy (rational):

but i feel it is a typo

OpenStudy (anonymous):

As do I

OpenStudy (anonymous):

I'm going to email my TA about it and see if she has had any other emails about the same problem. Thanks for the help

OpenStudy (rational):

good luck!